\begin{thebibliography}{9}

\bibitem{Archer} A. Archer \emph{A Modern Treatment of the 15 Puzzle.} American Mathematical Monthly 106 (1999). 793–799.
 
\bibitem{Conwaysurvey} J.H. Conway \emph{$M_{13}$.} London Mathematical Society Lecture Note Series, 241, pp. 1–11. Cambridge: Cambridge University Press, 1997. 

\bibitem{atlas}  J.H. Conway, R.T. Curtis, S.P. Norton, R.A. Parker and R.A. Wilson \textit{Atlas of finite groups. Maximal subgroups and ordinary characters for simple groups. With computational assistance from J. G. Thackray.} Oxford University Press, Eynsham, 1985.

\bibitem{CEM} J.H. Conway, N.D. Elkies and J.L. Martin \emph{The Mathieu group $\mathrm{M}_{12}$ and its pseudogroup extension $\mathrm{M}_{13}$.} Experiment. Math. 15 (2006), no. 2, 223–236. 

\bibitem{CS} J.H. Conway and N.J.A. Sloane \textit{Sphere packings, lattices and groups.} Third edition. With additional contributions by E. Bannai, R. E. Borcherds, J. Leech, S. P. Norton, A. M. Odlyzko, R. A. Parker, L. Queen and B. B. Venkov. Grundlehren der Mathematischen Wissenschaften [Fundamental Principles of Mathematical Sciences], 290. Springer-Verlag, New York, 1999. 

\bibitem{EJS} O. Ekenta, H.G. Jang and J.A. Siehler \emph{Slide-and-swap permutation groups. }Involve 7 (2014), no. 1, 41–55. 

\bibitem{GGPS} N. Gill, N.I. Gillespie, C.E. Praeger and J. Semeraro \emph{Conway groupoids, regular two-graphs and supersimple designs. }Sém. Lothar. Combin. 79 ([2018--2023]), Art. B79b, 24 pp.

\bibitem{GGS} N. Gill, N.I. Gillespie and J. Semeraro \emph{Conway groupoids and completely transitive codes. } Combinatorica 38 (2018), no. 2, 399–442. 

\bibitem{Jordan} M.W. Liebeck and J. Saxl
\emph{Minimal degrees of primitive permutation groups, with an application to monodromy groups of covers of Riemann surfaces.}
Proc. London Math. Soc. (3)63(1991), no.2, 266–314. 


\bibitem{Martin} J.L. Martin \emph{The Mathieu Group $\mathrm{M}_{12}$ and Conway’s $\mathrm{M}_{13}$-Game.} Under
graduate thesis, Harvard University, 1996.

\bibitem{SS} J. Slocum and D. Sonneveld \emph{The $15$ Puzzle Book: How it Drove the World Crazy. } Indiana University Press, 2006.

\bibitem{Wilson} R.M. Wilson \emph{Graph puzzles, homotopy, and the alternating group. } J. Combinatorial Theory Ser. B 16 (1974), 86–96.


  \end{thebibliography}

\documentclass[12pt]{article}
\usepackage{amsmath, amsthm, amssymb,amsfonts}
 \usepackage[utf8]{inputenc}
\usepackage{mathtools}
\makeatother
\usepackage{enumerate}
\usepackage{enumitem}  
\usepackage{hyperref}
\usepackage[margin=1in]{geometry}
\newtheorem{theorem}{Theorem}[section]
\newtheorem{thm}{Theorem}

\newtheorem{lemma}[theorem]{Lemma}
\theoremstyle{definition}
\newtheorem{remark}[theorem]{Remark}
\newcommand{\Stab}{\mathrm{Stab}}

\newcommand{\PGL}{\mathrm{PGL}}
\newcommand{\Sym}{\mathrm{Sym}}
\newcommand{\Alt}{\mathrm{Alt}}
\newcommand{\M}{\mathrm{M}}
\newcommand{\GF}[1]{\mathbb{GF}(#1)}
\newcommand{\PP}[1]{\mathbb{P}(#1)}
\renewcommand{\P}[1]{\mathcal{P}(#1)}
\renewcommand{\L}[1]{\mathcal{L}(#1)}
\renewcommand{\t}[1]{\tau_{#1}}
\newcommand{\h}[1]{h_{({#1})}}
\newcommand{\x}[1]{x_{({#1})}}
\newcommand{\Pseudo}[1]{\frak{G}(#1)}
\newcommand{\PseudoG}[1]{\mathcal{G}(#1)}
\begin{document}
\title{\Huge{ $\PP{q}$-groupoids of Conway type}}
\author{Veronica Kelsey\footnote{Email address of corresponding author: veronica.kelsey@manchester.ac.uk  \newline Key words: groupoids, projective planes \newline MSC: 20B25}, Peter Rowley \\
 \small{Department of Mathematics, University of Manchester, Oxford Road, M13 6PL, UK}\\ \small{Heilbronn Institute for
Mathematical Research, Bristol, UK}}
 \maketitle
\begin{abstract}
In the spirit of Conway \cite{Conwaysurvey} we define a groupoid starting from projective planes of order $q$, where $q$ is odd. The associated group of these groupoids is then investigated.

\end{abstract}
\date{}

\section{Introduction}\label{Sec:Intro}
Noyes Palmer Chapman's classic 15-Puzzle, dating back to the 1870s,  consists of 15 numbered tiles and one hole in a $4 \times 4$ tray. In this puzzle, often incorrectly credited to Sam Loyd (see Slocum and Sonneveld \cite{SS}), tiles are slid horizontally or vertically and the goal is to arrange the tiles in numerical order.

Taking inspiration from the 15-Puzzle Conway \cite{Conwaysurvey} proposed a variation based on replacing the $4 \times 4$ tray with the 13 points of the projective plane of order 3. Following on from Conway's idea there have been a number of other variations on this theme involving certain graphs or block designs. For an example of the former using regular $3$-valency graphs see Wilson \cite{Wilson} and Archer \cite{Archer} and for the latter, employing $2-(n,4,\lambda)$ designs, consult Gill, Gillesie, Praeger, Semeraro \cite{GGPS} and Gill, Gillesie, Semeraro \cite{GGS}.  Yet other perspectives may be seen in Ekenta,  Jang, Siehler \cite{EJS}. Here we study a further extension of these ideas but continuing with the theme of projective planes.\medskip 

We also emulate the 15-Puzzle construction. In place of a $4 \times 4$ tray, we consider the points $\P{q}$ of the projective plane $\PP{q}$. For each $\alpha,\beta \in \P{q}$ we will define a unique move $\h{\alpha,\beta}$ which interchanges $\alpha$ and $\beta$, acts as an element of order 2 on the line $\langle \alpha, \beta \rangle$, and fixes the remaining points of $\P{q}$.
Let $\Pseudo{q}$ be the set of elements given by sequences of elementary moves of the form 
\[\h{\alpha_0,\alpha_1,\ldots, \alpha_s}:=\h{\alpha_0,\alpha_1}\h{\alpha_1,\alpha_2}\cdots \h{\alpha_{s-1},\alpha_s}.\]
Two elements $\h{\alpha_0,\ldots, \alpha_s}, \h{\beta_0,\ldots, \beta_t} \in \Pseudo{q}$ are multiplied by concatenation provided $\alpha_s=\beta_0$ to give $\h{\alpha_0,\ldots, \alpha_s, \beta_1,\ldots, \beta_t}$. Since not all possible products are defined, $\Pseudo{q}$ is a groupoid and not a group.

If we fix $\alpha \in \P{q}$, then we have a group $\PseudoG{q}$ defined by
\[\PseudoG{q}: = \{ \h{\alpha,\alpha_1 , \ldots, \alpha_{s-1}, \alpha} \in \Pseudo{q} \} \leq \Sym\left(\P{q} \backslash \{\alpha\}\right) .\] 
 
Taking $q = 3$ recovers Conway's construction \cite{Conwaysurvey} yielding, as he termed it, the pseudogroup $\M_{13}$, which in our notation is $\Pseudo{3}$. Conway shows that for any $\alpha \in \P{3}$, the group $\PseudoG{3}$ is isomorphic to $\M_{12}$, the Mathieu group of degree $12$. Further, in \cite{Conwaysurvey}, Conway also considers what he terms the doubling of $\M_{13}$ which ends up exposing the double cover of $\M_{12}$. In Conway, Elkies and Martin \cite{CEM} (see also Martin \cite{Martin}) there is an extended account of both $\M_{13}$ and its doubling.\medskip

In this paper we determine $\PseudoG{q}$.
\begin{thm}\label{main} Let $q>3$ be an odd prime power, $\alpha$ a fixed point of $\P{q}$ and set $\Omega = \P{q} \backslash \{\alpha\}$. Then
\[\PseudoG{q}= \begin{cases}
\Alt(\Omega) & \text{ if } q \equiv 3 \bmod 4,\\
\Sym(\Omega)& \text{ if }q \equiv 1 \bmod 4.
\end{cases}\]
\end{thm}

The sporadic group $\M_{12}$ has a history of having tendrils in many, and varied, combinatorial areas. See \cite{atlas} and \cite{CS} for a summary of some of these. In relation to $\PseudoG{q}$, yet again $\M_{12}$ stands as a kind of singularity.

\section{Preliminaries}
We begin with some general facts on projective plans.
Let $q$ be a power of an odd prime and $\PP{q}$ denote the projective plane defined over $\GF{q}$. Hence $\PP{q}$ consists of a point set $\P{q}$ and a line set $\L{q}$, each of size $q^2+q+1$.
It is well-known that every distinct pair of points $\beta,\gamma \in \P{q}$ lie on a unique line which we denote by $\langle \beta, \gamma \rangle$. Dually, any two distinct lines $\ell, \ell' \in \L{q}$ intersect in a unique point which we denote by $\ell \cap \ell'$. For a line $\ell\in \L{q}$, let $\Gamma_0(\ell)$ denote the set of points incident with $\ell$. We use $\{e_1,e_2,e_3\}$ to denote the standard basis vectors of $\GF{q}^3$.

\begin{lemma}\label{cyclicaction} 
Let $L =\PGL(3,q)$ act on $\PP{q}$, let $\beta, \gamma \in \P{q}$ be distinct, and let $\ell=\langle \beta,\gamma\rangle$. 
\begin{enumerate}[label=\rm{(\roman*)}]
\item \label{beta,ell} $L_{\beta,\ell}: = \Stab_L(\beta) \cap \Stab_L(\ell)$ has shape $q^3(q-1)^2$ and acts $2$-transitively on $\Gamma_0(\ell) \setminus \{ \beta \}$.
\item \label{beta,gamma}  $L_{\beta,\gamma}: = \Stab_L(\beta) \cap \Stab_L(\gamma)$ has shape $q^2:(q - 1)^2$ with the $q^2(q-1)$ fixing all of $\Gamma_0(\ell)$, and the induced action on $\Gamma_0(\ell) \setminus \{ \beta , \gamma\}$ is cyclic of order $q - 1$.
\end{enumerate}
\end{lemma}
\begin{proof} Since $L$ is 2-transitive on $\P{q}$ we may assume that $\beta=\langle e_1 \rangle$ and $\gamma=\langle e_2 \rangle$. 

The induced elements of $L_{\beta,\ell}$, and of $L_{\beta,\ell}$'s action on $\Gamma_0(\ell)\backslash \{\beta\}$ can be represented by right multiplication of the following matrices where $a,b ,c \in \GF{q}^*$ and $d,e,f,g \in \GF{q}$.
\[ \begin{pmatrix}
1 & 0 & 0\\
d & a & 0\\
e & f & b
\end{pmatrix} \quad \quad
\begin{pmatrix}
1 & 0 \\
g & c
\end{pmatrix}  \]

The induced elements of $L_{\beta,\gamma}$ are those of $L_{\beta,\ell}$ with $d=0$; of $L_{\beta,\gamma}$ fixing all of $\Gamma_0(\ell)$ are those of $L_{\beta,\ell}$ with $a=1$ and $d=0$; and the induced action of $L_{\beta,\ell}$ on $\Gamma_0(\ell)\backslash \{\beta, \gamma\}$ are those of $L_{\beta,\ell}$'s action on $\Gamma_0(\ell)\backslash \{\beta\}$ with $g=0$. \end{proof}

Let $\t{\beta, \gamma}$ denote an involution in $L_{\beta,\gamma}$, whose support is restricted to the points on $\ell \setminus \{ \beta, \gamma\}$. By Lemma \ref{cyclicaction} its action on $\Gamma_0(\ell) \backslash \{\beta,\gamma\}$ is well-defined and $\t{\beta,\gamma}$ will be the product of $(q-1)/2$ transpositions. In the case of $\beta=\langle e_1 \rangle$ and $\gamma=\langle e_2 \rangle$ the action of $\t{\beta,\gamma}$ on $\ell \backslash \{\beta, \gamma\}$ is as follows
\[\t{\beta,\gamma} : \langle ae_1 + be_2 \rangle \mapsto  \langle ae_1 + be_2 \rangle \begin{pmatrix}1 & 0 \\ 0 & -1 \end{pmatrix} = \langle ae_1-be_2 \rangle \;\;\; \text{ for } a,b \in \GF{q}^*.\]

We mirror the 15-Puzzle construction as follows. Place numbered counters on $q^2+q$ of the points and leave the remaining point empty.  Following \cite{Conwaysurvey} we call this empty point the \emph{hole}. 
For each choice of counter $t$ we can define an elementary move. Let $\beta \in \P{q}$ be the hole and $\gamma \in \P{q}$ be the point covered by $t$. Then the corresponding \emph{elementary move}, $\h{\beta,\gamma}$, is given by moving the counter $t$ to $\beta$ (thus leaving $\gamma$ empty), and applying $\t{\beta,\gamma}$. 
It is then well defined to apply $\h{\delta,\epsilon}$ after $\h{\beta,\gamma}$ if and only if $\gamma=\delta$.

As introduced in Section \ref{Sec:Intro}, the groupoid $\Pseudo{q}$ consists of well defined concatenations of elementary moves. That is, the elements $\h{\beta_0,\beta_1,\ldots, \beta_s}:=\h{\beta_0,\beta_1}\h{\beta_1,\beta_2}\cdots \h{\beta_{s-1},\beta_s}$. The sequence $\beta_0,\beta_1,\ldots, \beta_s$ gives the path that the hole takes through $\P{q}$.
\section{Proof of Theorem \ref{main}}
Throughout this section, fix $q>3$ an odd prime power and let $G:=\PseudoG{q}$.
Recall that $\alpha$ is a fixed point of $\P{q}$. 
For $\beta, \gamma \in \Omega$ with $\beta \ne \gamma$, put
$$\x{\beta, \gamma} = \h{\alpha, \beta}\h{\beta, \gamma}\h{\gamma,\alpha}.$$
Observe that $G$ is generated by these $\x{\beta,\gamma}$.
We begin by proving a lemma about the elements $\x{\beta,\gamma}$ where $\alpha, \beta$ and $\gamma$ do not all lie on a common line of $\PP{q}$.

\begin{lemma}\label{notcollinear} Suppose that $\beta, \gamma \in \Omega$ are distinct, and that $\alpha, \beta, \gamma$ are not collinear in $\mathbb{P}(q)$. Then 
\[\x{\beta, \gamma} = (\beta,\gamma)\t{\alpha,\beta}\t{\beta,\gamma}\t{\gamma,\alpha}\]
with $\x{\beta,\gamma}$ a product of $\frac{3q - 1}{2}$ pairwise disjoint transpositions.
\end{lemma}

\begin{proof} By definition 
\[\h{\alpha,\beta} = (\alpha,\beta)\t{\alpha,\beta}, \quad \h{\beta,\gamma} = (\beta,\gamma)\t{\beta,\gamma} \quad \text{and} \quad \h{\gamma,\alpha} = (\gamma,\alpha)\t{\gamma,\alpha}\] where $\t{\alpha,\beta}, \t{\beta,\gamma}$ and $\t{\gamma,\alpha}$ are each products of $\frac{q-1}{2}$ pairwise disjoint transpositions. By assumption $\alpha,\beta$ and $\gamma$ are not collinear, and so the supports of $\t{\alpha,\beta}$, $\t{\beta,\gamma}$ and $\t{\gamma, \alpha}$ are pairwise disjoint. In addition, $(\alpha,\beta), (\beta,\gamma)$ and $(\gamma,\alpha)$ all commute with each of $\t{\alpha,\beta}, \t{\beta,\gamma}$ and $\t{\gamma,\alpha}$. Hence from $(\alpha,\beta)(\beta,\gamma)(\gamma,\alpha)  = (\alpha)(\beta,\gamma)$, we obtain the required expression for $\x{\beta,\gamma}$. The number of transpositions is $3(\frac{q-1}{2} ) + 1 = \frac{3q - 1}{2}$.
\end{proof}

We now consider the case of $\alpha, \beta$ and $\gamma$ being collinear. We begin with some notation.
For a line $\ell\in \L{q}$ define the following subgroup of $G$,
\[G(\ell) = \left\langle \x{\beta,\gamma} \;| \; \beta, \gamma \in \Gamma_0(\ell) \setminus \{ \alpha \}, \; \beta \neq \gamma \right\rangle.\]
Recall that $L = \PGL_3(q)$ and $L_{\alpha,\ell} = \Stab_L(\alpha) \cap \Stab_L(\ell)$.

\begin{lemma}\label{collinear} 
Suppose that $\beta, \gamma \in \Omega$ are distinct, and that $\alpha, \beta$ and $ \gamma$ are collinear in $\mathbb{P}(q)$. Let $\ell$ be the line they all lie on and let $g \in L_{\alpha,\ell}$. Then the following hold.
\begin{enumerate}[label=\rm{(\roman*)}]
\item $g^{-1}\x{\beta, \gamma}g = \x{\beta^g,\gamma^g}$
\item $\x{\beta, \gamma} \neq 1$
\item $L_{\alpha,\ell}$ normalizes $G(\ell)$
\item $G(\ell)$ acts transitively on $\Gamma_0(\ell) \setminus \{\alpha\}$
\end{enumerate}
\end{lemma}

\begin{proof} 

\begin{enumerate}[label=\rm{(\roman*)}]
\item Since $\t{\beta, \gamma}^g \in (\Stab_{L_{\alpha,\ell}}(\beta) \cap \Stab_{L_{\alpha,\ell}}(\gamma))^g= \Stab_{L_{\alpha,\ell}}(\beta^g) \cap \Stab_{L_{\alpha,\ell}}(\gamma^g)$ and $\t{\beta,\gamma}^g$ induces the same action as $\t{\beta^g,\gamma^g}$ on $\Gamma_0(\ell) \setminus \{ \beta^g, \gamma^g \}$, it follows that $\t{\beta,\gamma}^g
= \t{\beta^g,\gamma^g}.$ Hence 
\[\h{\beta,\gamma}^g =((\beta,\gamma)\t{\beta,\gamma})^g = (\beta^g,\gamma^g)\t{\beta^g,\gamma^g} = \h{\beta^g,\gamma^g}.\] 
Since $g$ fixes $\alpha$, it follows as above that $\h{\alpha,\beta}^g = \h{\alpha,\beta^g}$ and $\h{\gamma,\alpha}^g = \h{\gamma^g,\alpha}$. Therefore the result follows.

\item Fix $\alpha=\langle e_1+e_2 \rangle$. By part (i) we may assume that $\beta = \langle e_1 \rangle$ and $\gamma=\langle e_2 \rangle$. For $a,b \in \GF{q}$ we have the following.
\begin{align*}
\langle ae_1+be_2 \rangle^{\h{\alpha,\beta}} &= \begin{cases} 
\langle e_1 \rangle & \text{ if } a=b^{-1},\\
\langle e_1+e_2 \rangle & \text{ if } b=0, \\
\langle (a-2b)e_1-be_2 \rangle & \text{ otherwise.}
\end{cases}\\ \\
\langle ae_1+be_2 \rangle^{\h{\beta,\gamma}} &= \begin{cases} 
\langle e_2 \rangle & \text{ if } b=0, \\
\langle e_1 \rangle & \text{ if } a=0,\\
\langle ae_1-be_2 \rangle \quad\quad\quad\quad & \text{ otherwise.}
\end{cases} \\ \\
\langle ae_1+be_2 \rangle^{\h{\gamma,\alpha}} &= \begin{cases} 
\langle e_1+e_2 \rangle & \text{ if } a=0, \\
\langle e_2 \rangle & \text{ if } a=b^{-1},\\
\langle ae_1+(2a-b)e_2 \rangle & \text{ otherwise.}
\end{cases} 
\end{align*}
First assume that $q$ is not a 3-power. Then
\[\langle 2e_1+e_2 \rangle^{\h{\alpha,\beta}\h{\beta,\gamma}\h{\gamma,\alpha}}=\langle e_2 \rangle^{\h{\beta,\gamma}\h{\gamma,\alpha}}= \langle e_1 \rangle^{\h{\gamma,\alpha}}= \langle e_1+2e_2\rangle.\]
Hence $\x{\beta,\gamma}$ maps $\langle 2e_1+e_2 \rangle$ to $\langle e_1+2e_2 \rangle=\langle 2e_1+4e_2\rangle$. Since $q$ is not a 3-power, it follows that $\langle 2e_1+e_2 \rangle \neq \langle 2e_1+4e_2\rangle$. Therefore $\x{\beta,\gamma} \neq 1$

Next assume that $q=3^r$ for some $r>1$. Let $\omega$ be a primitive root of unity in $\GF{q}$, and so in particular $\omega \neq -1$. Then
\begin{align*} 
\langle e_1+\omega e_2 \rangle^{\h{\alpha,\beta}\h{\beta,\gamma}\h{\gamma,\alpha}}&=\left\langle (1+\omega)e_1-\omega e_2 \right\rangle^{\h{\beta,\gamma}\h{\gamma,\alpha}}\\
&= \langle (1+\omega)e_1+\omega e_2 \rangle^{\h{\gamma,\alpha}}\\
&= \langle (1+\omega)e_1+(-1+\omega) e_2 \rangle.
\end{align*}
Hence $\x{\beta,\gamma}$ maps $\langle e_1+\omega e_2 \rangle=\langle (1+\omega)e_1+(\omega+\omega^2)e_2\rangle$ to $\langle (1+\omega)e_1+(-1+\omega) e_2 \rangle$. Since $q \geq 9$, we have that $w^{2} \neq -1$ and so $\langle (1+\omega)e_1+(\omega+\omega^2)e_2\rangle \neq \langle (1+\omega)e_1+(-1+\omega) e_2 \rangle$. Therefore, it this case we also have $\x{\beta,\gamma} \neq 1.$

\item It follows by part (i) that the generating set for $G(\ell)$ is invariant under conjugation by $L_{\alpha,\ell}$.

\item By Lemma \ref{cyclicaction}\ref{beta,ell}, $L_{\alpha,\ell}$ is 2-transitive on $\Gamma_0(\ell) \setminus \{\alpha\}$, and so $L_{\alpha,\ell}G(\ell)$ is also. Recall that 2-transitive groups are primitive, and non-trivial normal subgroups of primitive groups are transitive. Hence $G(\ell) \trianglelefteq L_{\alpha,\ell}G(\ell)$ is transitive on $\Gamma_0(\ell)\backslash \{\alpha\}$. 
\qedhere
\end{enumerate}
\end{proof}

\begin{remark}
\begin{enumerate}[label=\rm{(\roman*)}]
\item If instead $q=3$, Lemma \ref{collinear}(ii) does not hold. In a certain sense, this allows for the emergence of $\M_{12}$. 
\item The elements $\x{\beta,\gamma}$ in Lemma \ref{collinear} have a surprising range of possible cycle types on $\Gamma_0(\ell) \setminus \{\alpha \}$. Here is a sample of the cycle types for some small values of $q$.

\begin{tabular}{|c||c|c|c|c|c|c|c|c|c|c|c|}
\hline
$q$ & 5 & 7 & 9 & 11 & 13 & 17 & 19 & 23 & 25 & 27 & 29\\ \hline
cycle type & $1^1.4^1$ & $7^1$ & $1^5.4^1$ & $1^2.3^3$ & $2^3.7^1$ & $3^3.8^1$ & $1^2.17^1$ & $23^1$ & $1^1.4^1.5^4$ & $1^3.4^6$ & $1^2.13^1.14^1$ \\ \hline
\end{tabular}
\end{enumerate}
\end{remark}

\begin{lemma}\label{primitive} $G$ acts primitively on $\Omega$.
\end{lemma}
\begin{proof}  For $\beta \in \Omega$, let $\ell_{\beta}=\langle \alpha, \beta \rangle$ and $\Gamma_{\beta}=\Gamma_0(\ell_{\beta}) \cap \Omega = \Gamma_0(\ell_{\beta})\backslash \{\alpha\}$.

We claim that if $G$ is imprimitive, then there exists a non-trivial block $B$ and a point $\beta \in B$ with $\Gamma_{\beta} \subseteq B \not\subseteq \Gamma_{\beta}$, from which we will derive a contradiction. \medskip

First we find a non-trivial block $B$ containing a point $\beta$ such that $B \not\subseteq \Gamma_{\beta}$.
If this holds for all blocks, then we are done. Hence assume that $C$ is a non-trivial block containing a point $\gamma$ such that $C \not\subseteq \Gamma_{\gamma}$. Then there exists $\delta \in C \backslash \{\gamma\} \subseteq \Gamma_{\gamma}$ and $\epsilon \in \Omega \backslash\Gamma_{\gamma}$. 
Now $C^{\x{\gamma,\epsilon}}$ contains $\delta^{\x{\gamma,\epsilon}} \in \Gamma_{\gamma}$ and $\gamma^{\x{\gamma,\epsilon}} = \epsilon \notin \Gamma_{\gamma}$. Hence let $B=C^{\x{\gamma,\epsilon}}$ and $\beta = \delta^{\x{\gamma,\epsilon}}$.\medskip

We now show $\Gamma_{\beta} \subseteq B$. Let $\zeta \in \Gamma_B$, by the above we may let $\eta \in B \backslash \Gamma_B$. By Lemma \ref{collinear}(iv) there exists $g \in G(\ell_{\beta})$ with $\beta^g=\zeta$. Hence $\zeta = \eta^{\x{\beta,\eta}g} \in B^{\x{\beta,\eta}g}$. Since $G(\ell_{\beta})$ fixes $\Omega\backslash \Gamma_{\beta}$ pointwise, it follows that $\beta^{\x{\beta,\eta}g}=\eta \in B \cap B^{\x{\beta,\eta}g}$. Hence $\zeta\in B$.\medskip

Hence the claim holds. Let $\theta \in \Omega \backslash B$. Since $\Gamma_{\beta} \subseteq B$, it follows that $\ell_{\theta} \neq \ell_{\beta}$, and so $\x{\beta,\theta}$ interchanges $\beta$ and $\theta$ and fixes $\Gamma_{\beta} \backslash \{\beta\}$ setwise. Therefore $\theta  \in B^{\x{\beta,\theta}}\backslash B$ and $\Gamma_{\beta} \backslash \{\beta\} \subseteq B \cap B^{\x{\beta,\theta}}$, a contradiction.
\end{proof}

We can now complete the proof of Theorem \ref{main}.

\begin{proof}[Proof of Theorem \ref{main}]
By Lemma \ref{primitive} $G$ acts primitively on $\Omega$. Let $\beta,\gamma \in \Omega$ be distinct and such that $\alpha,\beta, \gamma$ are collinear. Then $\x{\beta,\gamma}$ is an element of $G$ with support size at most $q+1$. It is a simple exercise to show that $q+1<2(\sqrt{|\Omega|}-1)$, and so $\Alt(n) \leq G$ by \cite[Corollary 3]{Jordan}. Since $\h{\beta,\gamma}$ is a product of $\frac{q + 1}{2}$ transpositions, $\x{\beta,\gamma}$ is an even permutation if and only if $\frac{q + 1}{2}$ is even, which is equivalent to $q \equiv 3 \bmod 4$ and completes the proof of the theorem.
\end{proof}

\end{document}